OpenStudy (anonymous):

an isoceles triangle has a base of length 2√3 m and sides of length 2 m.Determine the exact value of the apex angle, in radians

OpenStudy (anonymous):

Start by dividing it into 2 right triangles with a line starting from the tip.

OpenStudy (alexwee123):

geometry or trig? @Sukhvir

OpenStudy (anonymous):

trig

OpenStudy (anonymous):

If you divide it into 2 right triangles, then you get a halved angle and a halved base. You can then get the angle via arcsin, then double the angle.

OpenStudy (alexwee123):

 now use law of cosine... \[(2\sqrt{3})^{2}=2^{2}+2^{2}-2(2)(2)\cos(\theta)\]